Pablo Galindo Facultad de Mateméticas Departamento de Análisis Universidad de Valencia Spain E-mail address: Pablo.Galindo@uv.es

Polynomials generated by linear operators

Abstract: The class of Banach algebra valued n-homogeneous polynomials generated by the n^{th} powers of linear operators is the topic of this talk. We compare it with the finite type polynomials, as well as their closures in the norm topology. This a joint work with M. L. Lourenco and L. A. Moraes.